### 12 Feb 2024 v2

# Picasso: Memory-Efficient Graph Coloring Using Palettes With Applications in Quantum Computing

Abstract
A coloring of a graph is an assignment of colors to vertices such that no two neighboring vertices have the same color. The need for memory-efficient coloring algorithms is motivated by their application in computing clique partitions of graphs arising in quantum computations where the objective is to map a large set of Pauli strings into a compact set of unitaries. We present Picasso, a randomized memory-efficient iterative parallel graph coloring algorithm with theoretical sublinear space guarantees under practical assumptions. The parameters of our algorithm provide a trade-off between coloring quality and resource consumption. To assist the user, we also propose a machine learning model to predict the coloring algorithm’s parameters considering these trade-offs. We provide a sequential and a parallel implementation of the proposed algorithm.
We perform an experimental evaluation on a 64-core AMD CPU equipped with 512 GB of memory and an Nvidia A100 GPU with 40GB of memory. For a small dataset where existing coloring algorithms can be executed within the 512 GB memory budget, we show up to 68$\times$ memory savings. On massive datasets we demonstrate that GPU-accelerated Picasso can process inputs with 49.5$\times$ more Pauli strings (vertex set in our graph) and 2,478$\times$ more edges than state-of-the-art parallel approaches.
Index Terms:
Graph coloring, quantum computing, memory-efficient algorithms.
I Introduction
Given a graph $G(V,E)$, the problem of graph coloring is to assign a color to each vertex such that no two adjacent vertices are assigned the same color, while minimizing the number of colors used. Graph coloring is one of the central problems in combinatorial optimization with applications in various scientific domains [1]. Many algorithmic techniques have been developed to solve coloring in sequential, parallel, and distributed settings for many variants of the coloring problem [2]. All of these algorithms are memory-demanding since they require loading the entire graph and several auxiliary data structrues into the memory. For massive graphs, especially on limited-memory accelerators (GPUs), these algorithms easily exhaust the available memory, and the problem is exacerbated for dense graphs due to quadratic scaling of edges. Our work considers graphs that are $\approx 50\%$ dense ($|E|\approx|V|^{2}/2$) for which the current graph coloring approaches [3, 4, 5, 6] quickly run out of memory when processing large instances (§ VII).
Our quest for memory-optimized coloring techniques stems from an application of quantum algorithms to computational chemistry. Quantum computers have the potential to offer new insights into chemical phenomena that are not feasible with classical computers. To utilize quantum approaches effectively, it is indispensable to encode chemical Hamiltonians and chemical-inspired wave function ansätze (i.e., the assumptions of the wave function form) in a quantum-compatible representation, while adhering to the proposed theoretical framework. However, the direct encoding of chemical Hamiltonians and typical chemical-inspired ansätze, which usually grows as high-degree polynomials, is unscalable. This directly affects the efficiency and applicability of the corresponding quantum algorithms. For example, transforming the chemical Hamiltonian from the second quantization to spin operators yields Pauli strings (tensor products of 2 $\times$ 2 Pauli and identity matrices) that scale as $\mathcal{O}(N^{4})$, where $N$ is the number of basis functions spanning the Hamiltonian [7, 8, 9].
One near-term solution to improve the scaling of quantum algorithms is the unitary partitioning and its variants, which aim to find compact representations of a linear combination of Pauli strings (§ II). These strategies, predominantly based on the graph analysis of Pauli strings, typically yield a reduction ranging from 1/10 to 1/6 in the number of alternative unitaries for small test cases. The unitary partitioning problem reduces to a graph coloring problem, where the vertices of the graph represent the Pauli strings, and the edges represent whether these strings obey an anticommute relation (§ II). These graphs are large and dense. Therefore, the state-of-the-art approaches for the unitary partitioning problem can only solve small molecules with a few thousand Pauli strings, whereas the desired scale is on the order of $\mathcal{O}(10^{6}\sim 10^{12})$ of Pauli strings.
Building on recent developments in sublinear algorithms for graph coloring [10], we introduce a novel parameterized coloring algorithm, Picasso, where the parameters of the algorithm provide a trade-off between the quality of the solution (i.e., the number of colors used) and the resource consumption of the algorithm. We theoretically prove that Picasso has a sublinear space requirement (§ IV), making it attractive to implement on GPUs. We empirically show that Picasso can solve trillion edge graph problems arising from molecules with more than 2 million Pauli strings in under fifteen minutes. Many of the results we report are the first-of-its-kind for the corresponding molecule and basis set. Since our parametric algorithm considers the target number of colors, we show that high-quality coloring can be achieved by an aggressive choice of parameters (§ VII-A1). We design a machine learning prediction model to determine the parameters configuration to be used to achieve a given trade-off between coloring quality and resource requirements (§VI).
Although Picasso is designed to solve a specific problem in quantum computing, it can be used in a generalized graph setting where memory efficiency is needed. The main contributions of this work are:
  • We introduce a first-of-its-kind graph coloring algorithm Picasso to address the unitary partitioning problem in quantum computing with demonstrations for dense inputs with up to two million vertices and over a trillion edges.
  • We prove that, under a practical assumption Picasso has sublinear memory requirement with high probability.
  • We propose a machine learning approach to predict the configuration of the algorithm’s parameters that allows achieving a given trade-off between coloring quality, runtime and memory requirements
  • We demonstrate the practical efficiency of our approach. The GPU-accelerated Picasso enables to process inputs with 49.5×49.5$\times$49.5 × more Pauli strings and 2,478×2,478$\times$2 , 478 × more edges than existing state-of-the-art parallel approaches.
II Problem Formulation
Our work is driven by the challenge of identifying compact unitary representations of chemical Hamiltonians and strongly correlated wave functions to enable accurate and efficient quantum simulations. Typically, this challenge can be abstracted as determining how to solve a clique partitioning problem efficiently. In this section, we delve into the graph formulation of the problem.
Figure 1: An overview of the mapping problem solved as clique partition using graph coloring of the conflict graph for H2 molecule with sto-3g basis function.
II-A Quantum computing problem
In quantum simulations aimed at elucidating physical and chemical phenomena, both the wave function and the system Hamiltonian play crucial roles. The wave function encodes the probabilistic state of a quantum system, facilitating the computation of various physical properties, whereas the system Hamiltonian governs the temporal evolution of the system’s states according to quantum mechanical principles and represents the total energy of the system. In this context, the wave function is responsible for state preparation, and the system Hamiltonian controls the evolution of the state.
In these simulations, the processes of state preparation and evolution must be conducted through unitary operations. These are operations that maintain the norm (magnitude) of the state vector in Hilbert space, ensuring that the total probability remains one. To accomplish this, both the system Hamiltonian and the wave function generator must be reformulated as either a unitary operator or a linear combination of unitary operators. This reformulation is crucial because unitary operations are reversible and preserve the quantum information within the system, making them indispensable for the coherent manipulation of quantum states in simulations. In small-scale demonstrations, one could employ, for instance, the Jordan-Wigner, Bravyi-Kitaev, or parity techniques [11, 12], which rewrite the Hamiltonian and the wave function generator as a combination of Pauli strings, each being a unitary. A Pauli string represents a tensor product of a series of $2\times 2$ Pauli matrices, $\sigma_{x},\sigma_{y},\sigma_{z}$, and the $2\times 2$ identity matrix, $I$. However, these techniques encounter the curse of dimensionality in large-scale applications. For instance, in molecular applications, reformulating the molecular Hamiltonian represented in the $N$-spin-orbital basis set would require $\mathcal{O}(N^{4})$ Pauli strings. The number of the Pauli strings needed to reformulate the highly correlated wave function ansätze, such as the non-unitary coupled-cluster ansätz that include single and double excitations, would be even higher, i.e. $\mathcal{O}(N^{7\sim 8})$. The curse of dimensionality makes the subsequent quantum simulation, especially quantum measurements, extremely challenging. Compact unitary representation for both the Hamiltonians and the wave function generators addresses the curse of dimensionality and the associated quantum measurement overhead.
Mathematically, given a set of $n$ Pauli strings, $P=\{P_{1},P_{2},\ldots,P_{n}\}$ each of which is of length $N$, along with their coefficients $\{p_{1},p_{2},\ldots,p_{n}\}$, we aim to compute a smaller set of unitaries $\{U_{1},U_{2},\ldots,U_{c}\}$ with the corresponding coefficients $\{u_{1},u_{2},\dots,u_{c}\}$, such that:
$\sum_{i=1}^{c}u_{i}U_{i}=\sum_{j=1}^{n}p_{j}P_{j},~{}~{}c<n.$ (1)
II-B Connection to Clique partitioning and Graph coloring
A straightforward way for satisfying Eq. (1) is enabled through the following condition for $m$ Pauli strings [13]:
$\sum_{i,j=1,i\neq j}^{m}p_{i}^{*}p_{j}P_{i}P_{j}=0,\ \ m\geq 2,$ (2)
for which a necessary condition is enabled by the anticommute property between any two Pauli strings in the given set, i.e.,
$\{p_{i}P_{i},p_{j}P_{j}\}=p_{i}^{*}p_{j}P_{i}P_{j}+p_{j}^{*}p_{i}P_{j}P_{i}=0,~{}~{}i\neq j.$ (3)
The above relationships can be translated to a clique partitioning and coloring problem in a graph as follows. Given a set of Pauli strings, $P$, we generate a graph $G(P,E)$, where $E$ is the set of all possible (unique) pairs of Pauli strings in $P$ that anticommute, i.e., satisfy the Eq. (3). A clique (or a complete subgraph) of a graph is a set of vertices such that every (unique) pair of vertices in that set are connected with an edge in $G$. A clique in $G$ thus corresponds to satisfying Eq. (2). Our goal is to generate a set of cliques, as small as possible, in $G$, which forms a partition of $P$. We formally define the problem as follows.
Definition 1 (Clique partitioning)
Given a set of Pauli strings, $P=\{P_{1},P_{2},,\ldots,P_{n}\}$, and a graph $G(P,E)$ generated from $P$, the clique partitioning problem is to compute a collection of cliques, $U=\{U_{1},\ldots,U_{c}\}$, where $\bigcup_{i}U_{i}=P$, and $U_{i}\cap U_{j}=\emptyset$, for $i\neq j$ that minimizes the number of cliques, i.e., $c$.
The clique partitioning problem is NP-Complete [14], which makes any optimal polynomial time algorithm unlikely. This problem is equivalent to the well-known graph coloring problem, which requires finding the smallest set of colors such that each vertex of the graph is assigned exactly one color from the set, and no two endpoints of an edge have the same color. The clique partitioning of a graph $G$ is equivalent to the coloring of the complement graph $G^{\prime}$ [15]. Let $G^{\prime}(V,E^{\prime})$ be the complement of a graph $G(V,E)$, where the vertex set remains the same but the edge set, $E^{\prime}=\{\{V\times V\}\setminus E\}\setminus\{\{v,v\}:v\in V\}$. Since the vertices of $G^{\prime}$ with the same color must represent a clique in $G$, it is easy to verify that any proper coloring of $G^{\prime}$ provides a feasible solution to the clique partitioning of $G$. We further note that when $|E|\geq|E^{\prime}|$, computing clique partitioning via coloring of the complement graph is more efficient.
In this paper, we solve the unitary partitioning problem by formulating it as a clique partitioning problem, which in turn is solved using graph coloring on the complement graph. An illustration of the process is shown in Fig.1 using the clique partitioning to generate a compact representation of the H2 molecule with sto-3g basis function, with $N=4$, as an example. We begin with a set of Pauli strings as vertices and compute the edges of the graph using Eq.3. We then construct the complement graph and color it. Finally, we output the partition according to the color classes. In the example, 17 Pauli strings are shrunk to a set of 9 unitaries.
III Related Work
Mapping Pauli Strings: To facilitate efficient term-by-term measurement schemes in NISQ devices, efforts to minimize the number of terms representing the Hamiltonian and wave function generators have been essential. Previous developments in hybrid quantum-classical algorithms have primarily focused on classically grouping Pauli strings to reduce quantum measurements. Drawing inspiration from the concept of Mutually Unbiased Bases (MUB) in quantum information theory [16, 17]—which is associated with maximizing the information gained from a single measurement—early endeavors have leveraged the commutativity of Pauli strings. This includes qubit-wise commutativity [18], general commutativity [19, 9], and unitary partitioning [20], to define the edges when translating the grouping problem to a clique partitioning and coloring problem in a graph composed of these Pauli strings.
For general molecular cases explored in quantum simulations, these strategies have the potential to group $(4^{N_{q}}-1)$ $N_{q}$-qubit Pauli strings (excluding the identity string) into no more than $3^{N_{q}}$ groups, with further reductions to $\mathcal{O}(N_{q}^{2\sim 3})$ possible, albeit at the cost of introducing additional one/multi-qubit unitary transformations before measurement. However, the polynomial scaling inherent in these grouping schemes renders script-based large-scale applications unscalable. This necessitates the development of high-performance computing libraries specialized in graph analysis for this purpose.
Graph Coloring: Graph coloring has been studied extensively in literature. The application of graph coloring in automatic (or algorithmic) differentiation has led to the study of different types of graph coloring problems [2], and a software library of serial implementations called ColPack [3]. Sequential coloring algorithms are based on greedy methods, which, given an ordering of the vertices, employ the smallest feasible color for each vertex. In the worst case and for general graphs, all these ordering-based methods require $\Delta+1$ colors, where $\Delta$ is the maximum degree of the graph. The ordering methods include Largest Degree First (LF), Smallest Degree Last (SL), Dynamic Largest Degree First (DLF), and Incidence Degree (ID). While greedy coloring techniques provide reasonable quality in sequential settings, these algorithms have little to no concurrency in practice. For parallel settings, there are two primary algorithmic techniques: i) heuristics building on the idea of finding maximal independent sets, introduced in the pioneering work of [21, ] and extended by [22, ] (JP), ii) heuristics leveraging speculation, where parallel threads speculatively color vertices using the least available color. Conflicts resulting from concurrent execution are then corrected in an iterative manner [23]. Similar approaches have been extended to manycore or GPU implementations [24, 25, 5], and distributed-memory algorithms and implementations [26, 27]. In § VII we compare Picasso for quality (against  [3]) and performance (against  [25, 5]).
The existing parallel graph coloring algorithms, especially the single-node GPU solutions, fail to solve large graph problems due to memory limitations. The graphs need to be loaded into the GPU memory along with auxiliary data structures such as an array of “forbidden colors”, which can easily exhaust the available memory. The graphs generated by our target applications are large and dense. Thus, there exists a need for coloring algorithms that are memory efficient.
Graph Coloring in sublinear space: The recent seminal work of Assadi, Chen and Khanna [10] (ACK, henceforth), stduied the $(\Delta+1)$-coloring problem with sublinear space and time constraints. They developed the Pallete Sparsification Theorem, which reduces the $\Delta+1$-coloring of a graph $G(V,E)$ to a list-coloring problem in a subgraph of $G$ with only $\mathcal{O}(|V|\log^{2}{|V|})$ edges. Applications of this theorem is shown by designing algorithms in dynamic semi-streaming in a single pass, sublinear query and MPC models. In terms of semi-streaming algoirthm, the only known $(\Delta+1)$-coloring before ACK’s algorithm [10] was the $\mathcal{O}(\log{|V|})$ pass distributed algorithm of Luby [21, 28], simulated in streaming setup. In semi-streaming model, ACK uses palette size as $\Delta+1$, list size of colors as $\mathcal{O}(\log|V|)$ for each vertex, and a special post-processing step. In this paper, we put ACK’s algorithm in practice by non-trivial modifications listed as follows.
  1. i)
    The palette size in ACK’s algorithm is $\Delta+1$, which limits the practicality of the algorithm on problems operating on large dense graphs, like the one we considered. Our graphs (original and complement) are dense ($\Delta>|V|/2$). One would require a fraction of $|V|$ to color all the vertices, as shown in Table III in § VII-A1 for our dataset ($<=16$% of $|V|$). Our algorithm allows users to specify a variable palette size. Modifying the analysis of [10], we proved that if the ratio $\Delta/\mathcal{P}$ is bounded (by $\log{|V|}$), a sublinear space is guaranteed. This assumption holds for our graphs since this ratio is a constant for our use cases.
  2. ii)
    The conflict graph coloring algorithm of [10] decomposes the graph and then applies a greedy coloring, an almost clique coloring, and a maximum matching-based coloring on the decomposition. We provide an efficient implementation of list-greedy-based coloring that dynamically colors the vertices based on their color list size.
  3. iii)
    ACK’s streaming algorithm is single-pass. For valid coloring, the single iteration algorithm would require a large palette size, degrading the solution quality. We address this issue with an iterative approach, where in each iteration, we attempt to color the uncolored vertices from the previous iteration. Our proof for sublinear space holds for each iteration.
IV Our Algorithm
TABLE I: Notation used in the paper.
Symbol Description
$P$ Set of Pauli strings
$N$ Length of each Pauli string
$\ell$ Iteration counter
$G_{\ell}=\{V,E\}$ Complement graph from $V\subseteq P$ at it. $\ell$
$n$ number of vertices, $n\coloneqq|V|$
$G_{c}=\{V_{c},E_{c}\}$ Conflict graph, $V_{c}\subseteq V$
$\delta(v)$ Degree of vertex $v$, with neighbor set: $adj(v)$
$\mathcal{P}$ Palette size, Palette=$\{(\ell-1)\mathcal{P},1,\ldots,\ell\mathcal{P}-1\}$
$\mathcal{C}$ Final number of colors, $\mathcal{C}\leq\mathcal{P}$
$\alpha$ Multiplicative factor for List size
$\mathcal{L}$ List Size, set to $\alpha\log|V|$
$\beta$ Weighting factor for bi-objective optimization
Algorithm 1 Picasso: Palette-based graph coloring
1:Input: A graph $G=(V,E)$
2:Output: A $color$ array, with a valid coloring of $G$
3:Initialize the $color$ array
4:$\ell=1$triangleright The iteration number
5:$G_{\ell}\leftarrow$ $G$While $V$ is not empty
6:$(\mathcal{P}_{\ell},\mathcal{L}_{\ell})\leftarrow$ Initialize palette and list size for $G_{\ell}$
7:$colList\leftarrow$ assign_rand_list_colors($\mathcal{P}_{\ell},\mathcal{L}_{\ell},G_{\ell}$)
8:$G_{c}\leftarrow$ construct_conflict_graph(colList, $G_{\ell}$)
9:color_unconflicted_vertices($V$setminus $V_{c}$, $colList$, $colors$)
10:$V_{u}\leftarrow$ color_conflict_graph($G_{c}$, $colList$, $colors$)
$\ell=\ell+1$
12:$G_{\ell}\leftarrow$ subgraph induced by $V_{u}$ in $G$
13:$V\leftarrow V_{u}$EndWhile
Picasso, listed in Algorithm 1, attempts to color the graph by initially assigning to each vertex a list of candidate colors chosen uniformly at random from a palette of $\mathcal{P}$ colors $\{0,1,\ldots,\mathcal{P}-1\}$, and progressively assigning to each vertex a color from its color list without violating the graph coloring constraint. It takes a graph $G$ as input and computes a valid coloring of $G$ in the $color$ array. The algorithm starts with the original graph $G$, and iteratively computes coloring for a subgraph. In each iteration ($\ell$), it estimates the palette size ($\mathcal{P}_{\ell}$) and list size ($\mathcal{L}_{\ell})$ for the current subgraph $G_{\ell}$ (Line 6). Next, it assigns to each vertex a list of candidate colors chosen uniformly at random from the palette (Line 7). For each vertex, we record the list of colors in the $colList$ data structure, where $colList(u)$ refers to the list of colors assigned to vertex $u$. We build the conflict subgraph, which consists of the edges that share a common color in their corresponding lists (Line 8) using the $colList$ array. We then color the unconflicted vertices using an arbitrary color from their color list (Line 9), and attempt to color the conflict graph (Line 10). The vertices uncolored in the current iteration are denoted as $V_{u}$. We then compute the subgraph induced by $V_{u}$ and continue if $V_{u}$ is non-empty. We note that the algorithm in every iteration starts with a new palette of colors and attempts to color the subgraph with these new colors. The colors of an iteration are not reused in the subsequent iterations. We can ensure that by defining the palette set as $\{(\ell-1)\mathcal{P},\ldots\ell\mathcal{P}\}$ at iteration $\ell$.
We now describe the construction and coloring of the conflict graph at iteration $\ell$. For ease of presentation, we omit the subscript $\ell$.
IV-A Conflict Graph Construction
An edge $(u,v)$ in $G$ is conflicted if its two endpoints share a color, i.e., $colList(u)\cap colList(v)\neq\emptyset$. If the color values in $colList$ are sorted, we need $O(\mathcal{L})$ time to check for a conflict edge. In our application, we are not provided with the graph. Instead, we are given a set of Pauli strings $P$ that defines the vertex set of $G$. We use a bit encoding scheme to dynamically derive the edges from Pauli strings in a memory-efficient way. Given two Pauli strings $P_{i}$ and $P_{j}$, whether an edge exists between them (in our case, a non-edge) can be checked using Eq. 3. Here, a Pauli string consists of $N$ characters, where each character corresponds to a $2\times 2$ Pauli matrix:
$\sigma_{x}=\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right),\sigma_{y}=\left(\begin{array}[]{cc}0&-i\\ i&0\end{array}\right),\sigma_{z}=\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right).$ (10)
When directly implementing the anti-commute condition given by Eq. 3, we must carry out $N-1$ tensor products for each string, followed by two matrix multiplications for matrices $P_{i}$ and $P_{j}$. However, the inherent properties of Pauli matrices allow us to check this condition efficiently. Specifically, any pair of Pauli matrices will either commute or anti-commute. Importantly, two distinct Pauli matrices will anti-commute
$\{\sigma_{i},\sigma_{j}\}=0~{}~{}\text{if}~{}~{}\sigma_{i}\neq\sigma_{j}\neq I,$ (11)
which is simplified to an element-wise character comparison.
The anti-commutation property can be extended to Pauli strings. This is done by counting the element-wise character comparisons corresponding to the anti-commute relation. It is worth noting that the anti-commutation relationship between two Pauli strings is determined by phase ±$\pm i$. Only an odd number of element-wise anti-commuting character comparisons will result in a non-vanishing phase.
Further reduction in the number of comparisons can be achieved using bit representations. Specifically, we make an observation: as there are only four possible matrices ($\sigma_{x},\sigma_{y},\sigma_{z},I$), we can encode an 8-bit character datatype to a smaller representation using bits. If we encode each Pauli matrix into a 2-bit value, we would still need to perform $N$ comparisons to count the number of mismatches. To determine a complement edge, we only check whether the number of mismatches is even or odd, which is determined by observing the least significant bit of the count value. With this, we need an encoding that causes a bit flip only when there is a mismatch between $\sigma_{x},\sigma_{y},$ or $\sigma_{z}$.
To accomplish this, we implement an encoding scheme similar to an inverse one-hot encoding, where matrices $\sigma_{x},\sigma_{y},$ and $\sigma_{z}$ are assigned 110, 101, and 011, and $I$ is assigned 000, respectively. We perform a bitwise AND operation between the encoded Pauli strings and perform a popcount operation to count the number of ‘1’ bits. Due to this encoding scheme, the only time the least significant bit is flipped in the popcount is during a mismatch, allowing us to track whether there is an odd or an even number of mismatches. Speedups from the encoded implementation on CPU range from $1.4$ to 2.0×2.0$\times$2.0 ×, including the encoding overheads.
IV-B Coloring the Conflict Graph
Once we compute the conflict graph, $G_{c}$ we are required to color $G_{c}$ using the list of colors assigned to each vertex. Here, we discuss two possible approaches to achieve that.
Static order schemes: Given an ordered set of vertices, we iterate through the set and attempt to color each vertex with the first available color in its list that does not conflict with the colors already assigned to its neighbors. We may use popular [2] vertex order such as the Natural, Largest Degree First, Smallest Degree Last, or Random ordering.
Dynamic vertex order scheme: A second approach is to color the conflict graph using the lists in a dynamic order [29]. Here, we describe a dynamic greedy algorithm from [29] for the list coloring and discuss an efficient implementation of it. The algorithm attempts to color the vertices that are most constrained, i.e., it colors a vertex according to the size of the list. When a vertex is assigned a color from its list, the assigned color is removed from the list of all the neighbor of the vertex, rendering a dynamic order on the vertices. A naïve implementation of the dynamic list coloring algorithm would require $\mathcal{O}(|V_{c}|^{2}+|E_{c}|\mathcal{L})$ time. We have at most $|V_{c}|$ iterations, and in each iteration, we need to find the vertex with the smallest current list size in $\mathcal{O}(|V_{c}|)$ time, then color the vertex and mark this color removed from all its neighbors in $\mathcal{O}(\delta(v)\mathcal{L})$ time, where $\delta(v)$ is the degree of vertex $v$. Summing over $|V|$ gives us the total time. We can reduce it to $\mathcal{O}((|V_{c}|+|E_{c}|\mathcal{L})\log{|V_{c}|})$ using a minimum heap data structure with logarithmic update_key operation. Next, we present an efficient implementation in Algorithm 2 that eliminates the $\log{|V_{c}|}$ factor by using the bucketing technique.
Algorithm 2 Greedy list-coloring of conflict graph, $G_{c}$
1:Input: Conflict graph $G_{c}$, color list $colList(v)$, $\forall v\in V_{c}$.
2:Output: A coloring of $G_{c}$ using the lists, and the uncolored vertex set, $V_{u}$.
[…]
4:Insert $v$ to $B[colList(v).size()]$EndFor
5:Mark all vertices of $G_{c}$ as unprocessed
6:$V_{u}\leftarrow\emptyset$While $\exists$ an unprocessed vertex
7:Pick a vertex, $v$ from the lowest bucket
8:Color $v$ from $colList(v)$ chosen uniformly at random, say $c$
[…]
11:Mark $u$ as processed
12:$V_{u}=V_{u}\cup u$
[…]
14:Remove $u$ from its current bucket
[…]
Algorithm 2 stores the vertices of $G_{c}$ in an array of buckets $B$, according to the size of their color lists. The bucket at $B[i]$ holds the vertices $v\in V$, whose size of the $colList(v)$ is $i$. We define the lowest bucket from the array as the non-empty bucket with the smallest index. The algorithm marks all the vertices as unprocessed and continue until all the vertices are processed. In each iteration, the algorithm finds the lowest bucket and chooses a vertex uniformly at random from it. This vertex is colored with an arbitrary color from its list and marked as processed. We say the color is $c$. Then, the algorithm scans all neighbors of this vertex in $G_{c}$ and removes $c$ if it exists in their color list. If any neighbor’s list becomes empty, we mark this neighboring vertex as processed. The runtime of Algorithm 2 is $\mathcal{O}((|V_{c}|+|E_{c}|)\mathcal{L})$, since there might be at most $|V_{c}|$ iterations and in each iteration we require $\mathcal{O}(\mathcal{L})$ time to find the lowest bucket. The removals of a vertex from a bucket (Lines 9 and 14) can be implemented in constant time by an auxiliary array that stores the location of the vertex in the bucket, while the removal of a color from $colList$ (Line 10) takes $\mathcal{O}(\mathcal{L})$. Thus, processing a chosen vertex requires $\mathcal{O}(\delta(v)\mathcal{L})$ time, and summation over all vertices gives the total time.
IV-C Analysis of the Algorithm
Building on the analysis in [10], we now show that, under reasonable assumptions on palette size, with a high probability our algorithm constructs in each iteration a conflict graph that is sublinear in the graph size. Following the previous section, we will omit the subscript $\ell$ since the results hold for any iteration. Recall that $G_{c}$ is the conflict subgraph computed from $G$ and $n\coloneqq|V|$. We will require the following concentration results.
Lemma 1 (Chernoff-Hoeffding bound [30])
Let $X_{1},\ldots,X_{m}$ be $m$ independent binary random variables such that $Pr(X_{i})=p_{i}$. Let $X=\sum_{i=1}^{m}X_{i}$ and $\mu=\mathbb{E}[X]$. Then the following holds for $0<\gamma\leq 1$:
$Pr\left[X\geq(1+\gamma)\mu\right]\leq e^{\left(-\frac{\mu\gamma^{2}}{3}\right)}$ (12)
Lemma 2
At iteration, $\ell$, let $G(V,E)$ be the graph to be colored with Palette size $\mathcal{P}$, and the color list size for each vertex $O(\log n)$. Let the average and maximum degree of $G$ be $\bar{d}$ and $\Delta$, respectively. Then, the total number of colors found by Algorithm 1 is $\sum_{\ell}\mathcal{P}_{\ell}$. At iteration $\ell$ of the Algorithm 1, the following relations hold:
  1.  2.1)
    Expected degree of vertex $v\in G_{c}$ is $\mathcal{O}(\frac{\delta(v)}{\mathcal{P}}\log^{2}{n})$.
  2.  2.2)
    Assuming $\frac{\Delta}{\mathcal{P}}=\mathcal{O}(\log n)$, the maximum degree and the maximum number of edges in $G_{c}$ is $\mathcal{O}(\log^{3}{n})$ and $\mathcal{O}(n\log^{3}{n})$ respectively with high probability.
  3.  2.3)
    Assuming $\frac{\bar{d}}{\mathcal{P}}=\mathcal{O}(\log n)$, the expected number of edges in $G_{c}$ is $\mathcal{O}(n\log^{3}n)$.
Proof:
Due to the design of the algorithm, the total number of colors used is: $\sum_{\ell}\mathcal{P}_{\ell}$.
  1. 1.
    Let us consider vertex $v$, and let $T$ be the size of the color list for $v$. Let $X_{v,u}$ be a binary random variable that takes $1$ if the edge $(v,u)$ is in the conflict graph $G_{c}$, and 00 otherwise. Also, let $X_{v}=\sum_{u\in adj(v)}X_{v,u}$. Let us fix the colors in the list of $v$ as $c_{1},c_{2},\ldots c_{T}$. The probability that at least one of these colors is shared with a neighbor, $u$, i.e., $Pr(X_{v,u}=1)$ is $\mathcal{O}(\frac{T}{\mathcal{P}}$). So the expected degree of $v$ is $\mathbb{E}[X_{v}]=\sum_{u\in adj(v)}\mathcal{O}(\frac{T}{\mathcal{P}})=\mathcal{O}(\frac{\delta(v)}{\mathcal{P}}\log^{2}{n})$, since $T=\mathcal{O}(\log^{2}n)$.
  2. 2.
    Since the maximum degree is $\Delta$, the maximum expected degree in $G$ is $\mathcal{O}(\frac{\Delta}{\mathcal{P}}\log^{2}{n})=\mathcal{O}(\log^{3}n)$ according to Lemma (2.1) and our assumption on the ratio $\frac{\delta(v)}{\mathcal{P}}$. We will now show the high probability concentration result. We note that all $X_{v,u}$s are independent of each other. So, using the Chernoff-Hoeffding bound (Lemma 1) and setting $\gamma=1$, the probability that $Pr[X_{v}\geq 2(\log^{3}n)]\leq e^{-(\log^{3}n)/3}\leq\mathcal{O}(n^{-\log^{2}n})$. So with high probability the maximum degree of $v$ is $\mathcal{O}(\log^{3}n)$ assuming the ratio $\frac{\Delta}{\mathcal{P}}=\mathcal{O}(\log n)$.
  3. 3.
    We can further improve the bound by using the average degree ($\bar{d}$) of $G$ rather than the maximum degree on our assumption. Let $Y$ be the random variable representing the sums of degrees of the graph at level $\ell$. From Lemma (2.1), the expected sums of degrees,
    $\mathbb{E}[Y]=\sum_{v\in n}X_{v}=$ $\sum_{v\in V}\mathcal{O}\left(\frac{\delta(v)}{\mathcal{P}}\log^{2}{n}\right)$
    $=\sum_{v\in V}\mathcal{O}\left(\frac{\bar{d}}{\mathcal{P}}\log^{2}{n}\right)=\mathcal{O}(n\log^{3}n).$
    This relationship follows since the sums of degrees of a graph can be replaced with sums of average degrees. The final equality follows from our assumption, $\frac{\bar{d}}{\mathcal{P}}=\mathcal{O}(\log n)$.
V Parallel GPU implementation
Algorithm 3 GPU Conflict Graph Construction
1:Input: Pauli strings, $V$ and their list of colors $colList$
2:Output: Conflict graph $G_{c}$ in CSR format
3:AvailMem = min($2|V|(|V|-1)$, MaxAvailGPUMem)
4:Allocate AvailMem on the GPU
5:$V_{edgecount},E_{coo}\leftarrow$ build_unordered_coo(colList, $V$)
[…]
The conflict subgraph construction on Line 8 of Algorithm 1 significantly dominates the execution time for our application. Because of this, we implemented a parallel GPU version to reduce the bottleneck. We note that, despite the original graph being dense, in almost all the cases the conflict graph is expected to be significantly sparse (details in §VII-A1). Although the pairwise comparisons of vertices are independent of each other, due to the unknown number of conflicting edges at runtime, we designed an implementation capable of generating the conflict graph in Compressed Sparse Row (CSR) format that is not only efficient in memory usage during construction but also enables contiguous access of memory for processing the conflict graph. We present this implementation in Algorithm 3. As a preprocessing (now shown in the Algorithm 3), we copy the input data on the GPU with a size of $N\mathcal{L}|V|/10$ 4-byte values for the encoded Pauli strings and their list of colors, and initialize $2|V|$ edge offset counters. We use 8 bytes for the counter if $|V|^{2}\geq 2^{32}$; otherwise, the offset counters are 4-bytes. All remaining available memory on the GPU, or the worst-case edgelist size of $2|V|(|V|-1)$, whichever is smaller, is then allocated to store the unordered edgelist, and the conflict graph generation kernel is launched on the GPU. For this kernel, each thread processes one of the $|V|(|V|-1)/2$ possible edges and, and inspect whether the edge is both a complement and conflicting. If so, the edge is written in the output and the respective edge offsets are incremented. After the kernel execution, we are left with an unordered edge list of size $|E_{c}|$.The edges of the complement graph are determined independently during conflict graph construction, and the complement graph does not need to be stored on the GPU memory. Since for CSR representation each edge is stored twice, if $|E_{c}|$ used less than half of the available GPU memory, then we generate the CSR output on the GPU. Otherwise, we read the unordered edge list and convert it to CSR on the host CPU instead. We attempted warp-level reduction on the offset values to condense the number of global atomic operations, but the number of conflicting edges was much smaller than the total number of possible edges ($\leq$5in most cases), so the overhead outweighed the benefits. To get an estimation for $|E_{c}|$ to preallocate memory on the GPU, we developed a machine learning based predictor, which we describe next (§VI).
VI Prediction of Palette Size
We employ a machine learning (ML) based methodology to predict the palette size $\mathcal{P}$ and $\alpha$ values to simultaneously minimize the number of final colors $\mathcal{C}$ and the number of conflicting edges $|E_{c}|$. As these two objectives are conflicting, we introduce $\beta$ as the weighting factor to determine the balance between minimizing $\mathcal{C}$ and $|E_{c}|$. The goal for this prediction is to find the optimal combination of $(\mathcal{P},\alpha)$ to minimize $(\beta\cdot\mathcal{C}+(1-\beta)\cdot|E_{c}|)$, i.e.,
$\min_{(\mathcal{P},\alpha)}(\beta\cdot\mathcal{C}+(1-\beta)\cdot|E_{c}|).$ (13)
We generate a dataset by varying values for percentile palette size, $\mathcal{P}^{\prime}=\frac{\mathcal{P}}{|V|}\times 100$ (as a percentage of the number of vertices $|V|$ in the complement graph $G^{\prime}$) and $\alpha$ to perform a grid search. We capture the $(\mathcal{P}^{\prime},\alpha)$ combinations that minimizes Eq. (13) for different values of $\beta$. Finally, this dataset trains regression models to predict the $(\mathcal{P}^{\prime},\alpha)$ for a given graph $G$ and $\beta$. We train the regressor with several molecules and test its accuracy on a new set of molecules. The methodology can be summarized as follows:
  • Step 1: For a given graph $G(V,E)$, perform sweeps on $(\mathcal{P}^{\prime},\alpha)$ and compute $(\mathcal{C},E_{c})$.
  • Step 2: For a particular value of $\beta$, compute objective in Eq. (13), and select the optimal choice of $(\mathcal{P}^{\prime},\alpha)_{opt}$.
  • Step 3: Run Step 2 for different values of $\beta$, and collect corresponding values of $(\mathcal{P}^{\prime},\alpha)_{opt}$ to construct the data-set for the graph $G$.
  • Step 4: Run Steps 1-3 for the graphs correspond to different molecules to construct the complete training set.
  • Step 5: Train the regressor model with $(\mathcal{P}^{\prime},\alpha)$ as outputs of the model for a given graph $G$ and $\beta$.
  • Step 6: After the model is trained, we provide a new graph and a particular choice of $\beta$, for which, the model predicts an optimal choice of $(\mathcal{P}^{\prime},\alpha)$ that optimizes for Eq. (13).
Model Training and Results: We generated a dataset for the molecules provided in Table II for percentile palette sizes $\mathcal{P}^{\prime}\in\{1\%,2.5\%,5\%....,20\%\}$ and $\alpha\in\{0.5,1.0,....,4.5\}$. We capture the $\langle\mathcal{P}^{\prime},\alpha\rangle$ combinations that minimized Eq. (13) for $\beta\in\{0.1,0.2,....,0.9\}$. From the dataset, we used the first five molecules for training and the last two for testing the regression analysis. We experimented with several linear (ridge, lasso) and nonlinear predictors (svm-kernel-rbf, decision trees, random forests) [31]. For given input of: ($\beta,\ V,\ E)$, the nonlinear regression models performed better in predicting the ideal $\langle\mathcal{P}^{\prime},\alpha\rangle$ combination. In particular, the random forest regressor provided the best performance with multiple iterations, with a mean absolute percentage error (MAPE) of 0.19 and an R-squared value of 0.88 over 100 iteration runs. We selected the number of trees (estimators) to be 100 and the maximum tree depth of 20. This methodology provides Picasso a means to predict optimal choice of palette and list sizes for an input with values for ($\beta,\ V,\ E)$, where $\beta$ provides the trade-off. We note that while our model is trained specifically for the dense input graphs used in this work, it can be extended to any family of inputs that can be well characterized with data as described above.
VII Experimental Evaluation
TABLE II: Molecule Dataset in our experiment
Molecule Name # of qubits # of Pauli terms # of edges
H6_3D_sto3g 12 8,721 19,178,632
H6_2D_sto3g 12 18,137 82,641,188
H6_1D_sto3g 12 19025 90,853,544
H4_2D_631g 16 22529 127,024,320
H4_3D_631g 16 34481 297,303,496
H4_1D_631g 16 42449 450,624,984
H4_2D_6311g 24 154641 5,979,614,600
H4_3D_6311g 24 245089 15,017,722,736
H8_2D_sto3g 16 271,489 18,513,622,112
H8_1D_sto3g 16 274,625 18,944,162,720
H4_1D_6311g 24 312817 24,464,823,272
H8_3D_sto3g 16 419,457 44,149,092,736
H6_3D_631g 24 554,713 77,027,619,060
H10_3D_sto3g 20 1,274,073 410,446,230,804
H6_2D_631g 24 2,027,273 1,028,164,570,684
H6_1D_631g 24 2,066,489 1,068,358,440,628
H10_2D_sto3g 20 2,093,345 1,108,417,973,696
H10_1D_sto3g 20 2,101,361 1,116,895,244,280
We evaluate the performance and quality of the solution obtained using Picasso to the current state-of-the-art approaches. For our evaluation, we used a machine equipped with a 64-core AMD EPYC 7742 CPU, 512GB host DDR memory, and an NVIDIA A100 GPU with 40GB of HBM memory. Table II lists the datasets we selected for our evaluation. These systems are chosen with a deliberate intent to encompass a broad spectrum of quantum scenarios. By picking $H_{n}$ molecular systems with varying values of $n=4,6,8,10$, we aim to ensure diversity in system size, ranging from simple to complex structures, enabling us to gauge the algorithm’s scalability and performance across different magnitudes. The incorporation of three spatial configurations for each $H_{n}$ molecular system, namely 1D, 2D, and 3D, introduces dimensional variability, shedding light on the tool’s capability to manage problems with different geometric complexities and symmetries. As the size of the $H_{n}$ system increases, so does the intricacy of electron-electron interactions and correlations. By integrating systems of different sizes, the goal is to critically assess the tool’s proficiency in addressing varying degrees of electron correlation, a pivotal aspect in quantum calculations. Moreover, this diverse selection, spanning across multiple system sizes and dimensions, serves as an effective stress test for Picasso. It not only offers insights into its performance benchmarks but also aids in pinpointing potential areas of enhancement, ensuring a comprehensive evaluation of its reliability and robustness. We classify our datasets into three categories: i) Small ($\leq 10$ Billion edges); ii) Medium ($\leq 1$ Trillion edges); and iii) Large ($>1$ Trillion edges). We evaluate our implementation by considering:
  • Performance: We present a detailed performance study of the execution time of Picasso and compare it against two of the state-of-the-art GPU implementations of the distance-1 coloring: i) Kokkos-EB: edge-based coloring included as part of the kokkos-kernels [4, 25, 32], and ii) ECL-GC-R: shortcutting and reduction based heuristics for JP-LDF [33, 5].
  • Quality: We assess the quality of coloring obtained by Picasso with respect to sequential greedy coloring implementations in ColPack [2, 3], and GPU implementations of Kokkos-EB, and ECL-GC-R. We also study the quality versus memory trade-offs in these implementations.
  • Parameter Sensitivity: We study the impact of palette size ($\mathcal{P}$) and color list size ($\alpha,\mathcal{L}$) on the final coloring, the number of conflicting edges, and the runtime of Picasso.
We limit our relative comparisons only to the small dataset due to limitations imposed by specific implementations. In fact, ColPack and Kokkos-EB run out of memory beyond the small datasets (Kokkos-EB also runs out of memory for the last instance of the small dataset). ECL-GC-R does not support a graph size larger than what a 32-bit integer type can represent (2 to 4B). All results presented here are averaged over five runs. Five different seeds for pseudo-random number generation are used for Picasso runs. For Picasso, we present the results using only the greedy list coloring of Algorithm 2 to color the conflict graph since it provided better coloring relative to the static ordering algorithms.
We note that Picasso does not require loading the entire graph into memory to color it. Instead, it computes the conflicting subgraph on-the-fly at every iteration, thus providing the memory improvement. This is fundamentally different from the previous state-of-the-art algorithms. In fact, ColPack, Kokkos-EB, and ECL-GC-R require loading of the entire graph into memory before coloring it. Therefore, we decided to be conservative in comparing the performance of the different implementations, and we include the conflict subgraph construction time for Picasso while we exclude the graph construction time for all the other approaches. Under these conservative settings, we show that Picasso is better or comparable to ECL-GC-R and within a factor 2×2$\times$2 × slower than Kokkos-EB. We also note that the explicit construction of a complement graph is expensive for large instances.
The two main parameters of Picasso are the size of palettes ($\mathcal{P}$) and the color list ($\mathcal{L}$). At any iteration of the algorithm with $V$ as the vertices considered, in our experiments, $\mathcal{P}$ represents the percentage of vertices, and $\mathcal{L}=\alpha\log{|V|}$, where $\alpha$ is the coefficient. We omit the subscript since, in each iteration, the same percentage value and $\alpha$ are used. We report the number of colors, running time in seconds, and maximum resident set size in GB for memory. Apart from these, we also define a few other metrics as follows.
  • Color percentage: percentage ratio between the number of colors to the number of vertices of the input, i.e., $\frac{\mathcal{C}}{|V|}*100$. It represents the percentage of shrinkage of the Pauli strings to the unitaries (impact on the application).
  • Maximum Conflicting Edge percentage: The percentage ratio between the maximum number of conflicting edges (across all iterations of Picasso) to the number of complement edges of the graph, i.e., $\frac{|E_{c}|}{|E|}*100$.
VII-A Quality and Memory Comparisons
VII-A1 Small Dataset
TABLE III: Quality comparisons of the algorithms. Results in bold are the best coloring. Norm.: $\mathcal{P}=12.5\%,\alpha=2$; Aggr.: $\mathcal{P}=3\%,\alpha=30$.
Problem ColPack Picasso Kokkos-EB ECL-GC
LF SL DLF ID Norm. Aggr.
H6_3D_sto3g 2479 902 901 952 1425.4 880.6 1040.17 943
H6_2D_sto3g 5389 1598 1580 1634 2901.5 †1587.4 1749.17 1596
H6_1D_sto3g 5771 1672 1601 1689 3036.1 †1650.4 1815.83 1642
H4_2D_631g 10049 1922 1694 1917 3579.8 1784.2 1772.67 1860
H4_3D_631g 15883 2729 2633 2668 5431.4 2606 2478.50 2596
H4_1D_631g 19412 3241 2943 3233 6538 3212.8 3426.17 3098
H4_2D_6311g 72493 8615 6944 8628 22463.8 8917.4 NA NA
Tables III and IV show the number of colors achieved and the memory requirements of the compared algorithms for the small dataset, respectively. We experimented with four ordering heuristics for sequential greedy coloring that are commonly considered in the literature: Largest First Degree (LF), Smallest Degree Last (SL), Dynamic Largest Degree First (DLF), and Incidence Degree (ID). The previous study on unitary partitions [20, 13] only considered LF greedy coloring algorithm. We refer you to the excellent survey of [2, ] for details on vertex orderings.
[…]
TABLE IV: Memory comparison of the algorithms: Maximum resident memory in GB
Problem ColPack Picasso
Norm. Aggr. Kokkos-EB ECL-GC-R
H6_3D_sto3g 0.38 0.08 0.23 1.19 0.30
H6_2D_sto3g 1.52 0.16 0.90 4.50 0.77
H6_1D_sto3g 1.68 0.17 0.97 4.93 0.83
H4_2D_631g 2.72 0.20 1.24 6.81 1.10
H4_3D_631g 5.66 0.31 3.10 15.69 2.37
H4_1D_631g 10.77 0.38 4.31 23.69 3.51
H4_2D_6311g 140.23 2.06 57.12 NA NA
Table IV shows the maximum resident memory in GB of the reference implementations for the small dataset. We find that the normal configuration of Picasso is the most memory efficient. In particular, it requires 68$\times$ lesser memory than the ColPack for H4_2D_6311g while Kokkos-EB and ECL-GC-R run out of memory and couldn’t compute a solution for the same instance. We observe between 14×14$\times$14 × and 60×60$\times$60 × lower memory utilization when comparing Picasso’s normal mode to Kokkos-EB, and a reduction in memory usage of 5$\approx$ 5$\times$ 5 × when considering Picasso’s aggressive mode. ECL-GC-R shows to be more memory efficient than Kokkos-EB and is comparable to Picasso’s aggressive mode. However, ECL-GC-R memory optimizations come with runtime penalties that we will detail in our performance evaluation (§ VII-B).
VII-A2 Medium and Large Dataset
[…]
Figure 2: Input dataset scaling on the iterative GPU implementation up to 2 million vertices. $\alpha=2$, $\mathcal{P}=12.5\%$. The black dashed line in the top plot denotes the maximum conflicting edge ratio supported by a 40GB NVIDIA A100.
[…]
VII-B Performance Evaluation
VII-B1 CPU-only Vs. GPU-assisted Implementation
We report the speedup of our GPU implementation over the CPU-only implementation of Picasso in Table V. We show the average time over five runs for the conflict graph construction and the total time of the CPU-only implementation in the $2^{nd}$ and $3^{rd}$ columns, respectively. Here, the reported conflict graph construction time includes the cumulative time spent in building the conflict graphs during each iteration of the algorithm. The process of building conflict graph accounts for over $98\%$ of execution time (geo. mean) for these problems. We accelerate the conflict graph build on GPUs (§ V). The last two columns of Table V report the speedup of our GPU implementation with respect to the conflict graph build step and the total runtime. We see that as the problem size increases, the speedup also increases, and we expect that trend to continue for even larger problems. We report results only for the small datasets because we used a cut-off time of 1 hour, and the CPU implementation was able to complete only the small dataset within that time budget. The geometric mean of the speed up for the conflict graph construction step is $\sim$60$\times$, and that results into a $\sim$16$\times$ speed up for the entire application (geo. mean). We note that our GPU implementation produces exactly the same coloring as the CPU-only one because the conflict graph construction is deterministic.
TABLE V: Runtime comparison for CPU only and GPU assisted implementation. $\mathcal{P}=12.5\%,\alpha=2$
CPU only GPU assisted
Problem Graph Build Total Graph Build Total
Time(s) Time(s) Speedup Speedup
H6_3D_sto3g 3.14 3.26 24.37 2.21
H6_2D_sto3g 14.87 15.19 43.76 8.94
H6_1D_sto3g 16.35 16.69 45.36 9.75
H4_2D_631g 24.47 24.92 51.55 13.41
H4_3D_631g 57.67 58.40 73.21 26.30
H4_1D_631g 91.02 92.00 83.85 35.25
H4_2D_6311g 1,428.94 1,436.10 173.36 110.90
Geo. Mean 59.54 15.98
VII-B2 Performance on Medium and Large Dataset
[…]
Figure 3: Input dataset scaling on the iterative GPU implementation up to 2 million vertices. $\alpha=2$, $\mathcal{P}=12.5\%$.
Fig. 3 shows the running time (with a breakdown in components) for all the medium and one of the large datasets for our GPU implementation. The problems are sorted from left to right, with the smallest problem on the left. For all the inputs, we set $\mathcal{P}=12.5\%$, and $\alpha=2$. We see that the conflict coloring that happens in CPU dominates the runtime. Despite this, we were able to color the largest graph with over 1 Trillion edges within 800 seconds with a color percentage ranging between 14–15%. For detailed quality results, see § VII-A2.
VII-C Performance Comparison with Kokkos-EB and ECL-GC-R
[…]
Figure 4: Performance comparison of Picasso and Kokkos-EB on the small datasets relative to ECL-GC-R execution time. For Picasso runs, $\mathcal{P}$ is varied and $\alpha=4.5$
Fig. 4 compares Picasso to the current state-of-the-art GPU implementations of graph coloring: Kokkos-EB and ECL-GC-R. The study is limited to the small dataset due to memory constraints imposed by specific implementations. We study quality of solution in terms of number of colors produced by the implementations, memory usage, and execution time. The results in Fig. 4 are normalized with respect to ECL-GC-R. We fixed $\alpha=4.5$ and we varied $\mathcal{P}$ from $1\%$ up to $15\%$ in our Picasso runs. We observe that the quality of solution achieved by Picasso increases when reducing the palette size ($\mathcal{P}$). When $\mathcal{P}=1\%$, Picasso matched the quality of solution of Kokkos-EB and ECL-GC-R (within $5\%$ – 15 %).
[…]
VII-D Parameter Sensitivity
[…]
Figure 5: Impact of $\mathcal{P}$ and $\alpha$ using H4_2D_6311g on final colors, number of conflicting edges and runtime for different inputs (lighter color is better).
The heatmap in Fig. 5 shows the impact of $\mathcal{P}$ and $\alpha$ on the final colors, conflicting edges, and execution time for a representative input (H4_2D_6311g). The heatmaps show normalized quantities with respect to what is observed for the same input graph. Therefore, as a measure of quality we report the fraction of colors with respect to the number of vertices in the input (lower is better) while we report execution time and the normalized fraction of the total number of conflicting edges processed by the algorithm as measures of the work done by the algorithm (lower is better).
The general trend favors smaller $\mathcal{P}$ and larger $\alpha$ to achieve a lower number of final colors at the cost of extra work. Conversely, larger $\mathcal{P}$ and smaller $\alpha$ lead to lower conflicting edges, and therefore, lower memory requirements and lead to faster execution time. We observed similar trends on all our datasets. These observations led us to design the approach in § VI to jointly minimize work and the size of the coloring by controlling $\mathcal{P}$ and $\alpha$.
VIII Conclusions and Future Work
We demonstrated the memory efficiency of Picasso using a large set of inputs and compared its performance with state-of-the-art approaches for graph coloring. In the realm of quantum computing, the introduction of Picasso marks a notable advancement. It is, quite plausibly, the inaugural scalable graph algorithm and implementation tailored for Hamiltonian and wave function partitioning that surpasses the capabilities of contemporary state-of-the-art quantum emulators. What amplifies its significance is its versatile nature; the same tool can be adeptly employed in qubit tapering, thereby reducing the effective number of qubits required for a given problem. When synergized with other methodologies, such as single reference guidance [13], Picasso promises the ability to solve systems comprising 100 to 1000 spin orbitals (i.e. qubits). It is within these vast and complex systems that the much-vaunted quantum advantage is believed to manifest, positioning our tool at the forefront of a transformative computational frontier.
Our future work will focus on developing distributed multi-GPU parallel implementations along with new algorithms for predicting $\mathcal{P}$ and $\alpha$ values to enable better trade-offs for quality and performance. We plan to further optimize our algorithm to address inputs from diverse applications with varying degrees of sparsity. We also plan to develop efficient algorithms for clique partitioning that explore applications beyond quantum computing. To the best of our knowledge, this is the first of its kind study for computing coloring of dense graphs on limited memory accelerator platforms and believe that it will enable several applications that critically depend on the computation of clique partitioning and graph coloring, as well as enable the development of memory-efficient graph algorithms.
Acknowledgement
The research is supported by the Laboratory Funded Research and Development at the Pacific Northwest National Laboratory (PNNL), the U.S. DOE Exascale Computing Project’s (ECP) (17-SC-20-SC) ExaGraph codesign center at PNNL, and NSF awards to North Carolina State University.